### v2

# Robustness of LiDAR-Based Pose Estimation: Evaluating and Improving Odometry and Localization Under Common Point Cloud Corruptions

Abstract
Accurate and reliable pose estimation, i.e., determining the precise position and orientation of autonomous robots and vehicles, is critical for tasks like navigation and mapping. LiDAR is a widely used sensor for pose estimation, with odometry and localization being two primary tasks. LiDAR odometry estimates the relative motion between consecutive scans, while LiDAR localization aligns real-time scans with a pre-recorded map to obtain a global pose. Although they have different objectives and application scenarios, both rely on point cloud registration as the underlying technique and face shared challenges of data corruption caused by adverse conditions (e.g., rain). While state-of-the-art (SOTA) pose estimation systems achieved high accuracy on clean data, their robustness to corrupted data remains unclear. In this work, we propose a framework to systematically evaluate five SOTA LiDAR pose estimation systems across 18 synthetic real-world point cloud corruptions. Our experiments reveal that odometry systems degrade significantly under specific corruptions, with relative position errors increasing from 0.5% to more than 80%, while localization systems remain highly robust. We further demonstrate that denoising techniques can effectively mitigate the adverse effects of noise-induced corruptions, and re-training learning-based systems with corrupted data significantly enhances the robustness against various corruption types.
I Introduction
Autonomous robots and vehicles rely on real-time pose estimation to navigate, avoid collisions, and perform tasks such as mapping and path planning. LiDAR enables precise environmental perception by generating high-resolution 3D point clouds independent of external lighting conditions. As a result, it is widely used for pose estimation, particularly in scenarios where GNSS signals are unreliable (e.g., urban canyons) or unavailable (e.g., tunnels). Two primary methods are commonly employed for LiDAR-based pose estimation: odometry [1] and localization [2]. LiDAR odometry estimates incremental motion by registering consecutive scans, making it essential for exploring unknown environments. However, odometry suffers from drift due to error accumulation over time. LiDAR localization, on the other hand, aligns online scans with pre-built maps to achieve drift-free global pose estimation. This is particularly crucial in safety-critical applications, such as fully autonomous vehicles that require global localization accuracy at the centimeter level [3]. Despite the different objectives, both tasks share underlying point cloud registration techniques to obtain fine poses.
Point cloud data corruptions can occur in various real-world scenarios, including adverse weather [4], noises from sensor [5] or dusts [6], reflective surfaces [7], and occlusions caused by dynamic environments [8, 9]. Despite SOTA LiDAR-based odometry and localization methods achieving highly accurate results on clean LiDAR datasets, corrupted point cloud data will inevitably bring challenges to their performance [10]. However, the extent to which their performance degrades under these corruptions remains unclear. Severe performance degradation could lead to critical failures, such as robots getting lost [11] and autonomous vehicles driving off-road [12]. Though prior studies explored the robustness [9] and security [13] of LiDAR-based localization systems, there is still a lack of systematic evaluation of how point cloud corruptions impact LiDAR-based pose estimation systems. Our work bridges this gap with three key contributions:
  • A framework to systematically evaluate the robustness of five SOTA LiDAR-based pose estimation systems under 18 common point cloud corruptions.
  • An investigation of mitigation strategies, including denoising and re-training (for learning-based systems).
  • Public release of code and data for reproducibility [^1] https://github.com/boyang9602/LiDARLocRobustness.
II Background and related works
II-A LiDAR-Based Pose Estimation
LiDAR-based pose estimation allows autonomous agents to track their motions (odometry) and get their absolute position and orientation in a map (localization) in 3D environments. The core of LiDAR-based pose estimation is point cloud registration, i.e., aligning LiDAR scans to estimate their 6-degree-of-freedom (DoF) spatial transformations (translation and rotation).
II-A1 Point Cloud Registration
Given a source point cloud $\mathcal{P}\in\mathbb{R}^{3\times N}$ and a target point cloud $\mathcal{T}\in\mathbb{R}^{3\times M}$, point cloud registration solves:
$\mathbf{R}^{*},\mathbf{t}^{*}=\underset{\mathbf{R}\in SO(3),\mathbf{t\in\mathbb{R}^{3}}}{\arg\min}\ \text{dist}\left(\mathbf{R}\mathcal{P}+\mathbf{t},\mathcal{T}\right)$ (1)
Here, $\mathbf{R}$ is a 3-DoF rotation matrix, and $\mathbf{t}$ is a 3-DoF translation vector. The operation $\mathbf{R}\mathcal{P}+\mathbf{t}$ rotates all points using $\mathbf{R}$ and translates them by $\mathbf{t}$. The function $\text{dist}(\cdot,\cdot)$ quantifies the errors between the corresponding points in $\mathbf{R}\mathcal{P}+\mathbf{t}$ and $\mathcal{T}$. In other words, this equation seeks the optimal pose $\mathbf{R}^{*},\mathbf{t}^{*}$ that aligns the source point cloud $\mathcal{P}$ to the target point cloud $\mathcal{T}$.
Various solutions have been proposed for point cloud registration, evolving from geometry-based methods to learning-based methods. They can be broadly categorized into direct matching, feature-based, and learning-based methods [1]. Direct matching methods, such as Iterative Closest Point (ICP) [14], align raw points by iteratively minimizing the distance between nearest neighbor points. While vanilla ICP struggles with problems such as sensitivity to initialization and noise, modern variants such as KISS-ICP [15] address these limitations and remain competitive currently. Another direct method, Normal Distribution Transform (NDT) [16], partitions the target point cloud into a grid of cells, with each cell modeling the local distribution of points as a Gaussian. Then it finds the optimal pose that transforms the source point cloud to best fit the normal distribution in the target grid. Feature-based methods like LiDAR Odometry and Mapping (LOAM) [17] achieve fast and robust odometry by extracting and matching feature points such as sharp edges and planar surfaces. With the development of deep learning, researchers further improve feature extraction from hand-crafted features to learned features and predict the pose using neural networks [18]. Instead of processing the 3D point cloud directly, [19] projects the points into 2D images and leverages the neural network to extract features from images and predict the pose. The success of Neural Radiance Fields (NeRF) [20] further inspires the recent exploration of implicit neural representations of scenes and advances the LiDAR-based pose estimations [21, 22].
II-A2 LiDAR Odometry
LiDAR odometry tracks the accurate motion of autonomous agents in real-time by aligning consecutive LiDAR scans. Despite the high accuracy of point cloud registration, the long-term trajectory drifts significantly due to accumulated errors. To mitigate the drift, additional algorithms such as loop closure [23] should be applied in post-processing. SOTA odometry systems cover all the approaches mentioned in the previous section [15, 17, 24, 19, 18, 22].
II-A3 LiDAR Localization
LiDAR localization determines the global 6-DoF pose by aligning an online scan with a pre-built map, which is critical for fully autonomous driving systems that require centimeter-level accuracy. While some localization systems estimate only 3-DoF position, we focus on 6-DoF pose estimation in this work. Compared to scan-to-scan registration in odometry, scan-to-map registration in localization is more challenging due to the vast search space. To address this problem, most methods first obtain an initial pose via external signals (e.g., GNSS) [3] or global position recognition [2], then perform the point cloud registration to estimate the fine pose. Although one-stage global pose estimation approaches exist, they generally underperform compared to those with initialization [2]. Similar to odometry, SOTA localization approaches also cover the full spectrum of registration methods [3, 25, 26, 21].
II-A4 Multi-Sensor Fusion Approaches
Multi-sensor fusion (MSF) combines complementary sensors, such as LiDAR, cameras (vision odometry and localization), IMUs (high-frequency motion data), and GNSS (global localization), to enhance pose estimation [3]. In this work, we focus on LiDAR-only pose estimation systems to study LiDAR’s standalone vulnerabilities in pose estimation.
II-B Robustness of LiDAR-based Systems
Previous research has extensively explored the robustness and security of LiDAR-based 3D perception systems, particularly in obstacle detection. Several studies have proposed realistic adverse weather simulations for point cloud, including rain [27, 28], snow [29], and fog [30]. Researchers also systematically benchmarked the LiDAR-based object detection on common point cloud data corruptions [31] and investigated its vulnerability to adversarial attacks [32].
In the domain of LiDAR-based pose estimation, recent works have examined the security of localization and odometry against adversarial attacks [13, 33, 34]. The most relevant study to ours is [9], which proposed a closed-form approximation of localization errors for the ICP algorithm. In contrast, our study systematically evaluates the robustness of multiple pose estimation systems, covering various underlying point cloud registration algorithms, against a comprehensive set of point cloud data corruptions.
III Methodology
This section presents our framework for evaluating and improving the robustness of LiDAR-based pose estimation.
Category Perturbation Type
Weather rain, snow, rain + wet ground (rain_wg),
snow + wet ground (snow_wg), fog
Noise background noise (bg_noise), upsample,
uniform noise in CCS (uni_noise),
gaussian noise in CCS (gau_noise),
impulse noise in CCS (imp_noise),
uniform noise in SCS (uni_noise_rad),
uniform noise in SCS (gau_noise_rad),
uniform noise in SCS (imp_noise_rad),
Density local density increase (local_inc),
local density decrease (local_dec),
beam deletion (beam_del),
layer deletion (layer_del), cutout
TABLE I: The 18 perturbations used by our work from three categories. CCS refers to Cartesian Coordinate System and SCS refers to Spherical Coordinate System. The details will be introduced in Section III.
III-A Evaluating Robustness of LiDAR-Based Pose Estimation
Autonomous robotics and vehicles are deployed in complex and diverse environments which affects the LiDAR performance and, consequently, the accuracy of pose estimation. To rigorously evaluate the robustness of pose estimation systems, our framework employs a broad spectrum of point cloud corruption (18 different types) that simulates adverse weather, internal and external noises, equipment malfunctions, and occlusions. Table I summarizes the list of corruption types.
Weather Perturbations. Adverse weather can negatively affect the quality of LiDAR scans. Randomly distributed droplets in the air from rain, snow, and fog can scatter light and disturb the measurement of LiDAR, and cause corruptions in point cloud. Furthermore, wet ground always co-occurs with rain and snow due to the precipitation, which reflects differently from dry ground. Despite the availability of real-world datasets capturing various adverse weather conditions [35, 36], we opt for simulation methods on clean datasets for the following reasons:
  • Realism. Existing simulation methods are physics model-based and highly resemble real-world phenomena [31, 27, 29, 30].
  • Controllable. In real-world datasets, the severity of adverse weather is typically unknown and cannot be adjusted. Simulation, on the other hand, allows us to parameterize the intensity of these conditions.
  • Sensor limitations. Real-world datasets are limited by their specific sensors, which may be incompatible with certain algorithms [10]. Simulation allows us to apply adverse weather on any dataset.
  • Completeness. No dataset contains the full range of corruptions. Simulation allows us to apply all corruptions on a single dataset for a comprehensive and fair comparison.
Figure 1: An example of rain and wet ground corruption. The field view and reflection intensities are significantly decayed compared to the clean data. The color bar indicates the intensities of the points.
Our framework simulates the following weather conditions considering both the direct impact of weather conditions on LiDAR (e.g., droplets causing scattering), as well as impacts of wet grounds on LiDAR (e.g., reflections).
  • Rain or Snow. We utilize an off-the-shelf weather simulator, i.e., LISA [27, 28] to simulate rain and snow. LISA utilizes a physics model to simulate the decayed return of laser pulses in rain and snow weather.
  • Fog. We utilize Fog Simulation (FS) [30] to simulate fog. FS models fog by modifying the impulse response changes to reflect the impact of different fog densities on the environment and the surrounding objects.
  • Rain or Snow + Wet Ground (rain/snow_wg). We combine LISA and Wet Ground Model from [29] to produce a more realistic simulation than rain/snow alone. The Wet Ground Model builds an optical model to simulate the reflection on the wet ground based on the thickness of the water. Figure 1 shows an example of this corruption.
Noise Perturbations. Noises occur in LiDAR measurement due to internal (e.g., equipment vibration) or external (e.g., dust) factors [37, 38]. We design perturbations on point cloud to simulate the errors attributed to such factors.
Noise Type Formula to Generate Noise
Gaussian (CCS) $p_{i}+\Delta p*SEV,\;\Delta p\sim\mathcal{N}(0,1)^{3},p_{i}\in P$
Uniform (CCS) $p_{i}+\Delta p*SEV,\;\Delta p\sim\mathcal{U}(-1,1)^{3},p_{i}\in P$
Impulse (CCS) $p_{i}+\Delta p*C,\;\Delta p\sim\mathcal{U}\{-1,1\}^{3},$
$p_{i}\in RS(P,SEV)$
Gaussian (SCS) $p_{i}+\Delta r*SEV,\;\Delta r\sim\mathcal{N}(0,1),p_{i}\in P$
Uniform (SCS) $p_{i}+\Delta r*SEV,\;\Delta r\sim\mathcal{U}(-1,1),p_{i}\in P$
Impulse (SCS) $p_{i}+\Delta r*C,\;\Delta r\sim\mathcal{U}\{-1,1\},$
$p_{i}\in RS(P,SEV)$
Background (CCS) $P_{new}=P\cup\{p_{i}^{\prime}|j=1,2,...,SEV\}$
$p_{i}^{\prime}=\{x_{i}^{\prime},y_{i}^{\prime},z_{i}^{\prime}\},\left\{\begin{array}[]{l}x_{i}^{\prime}\sim\mathcal{U}(x_{\min},x_{\max})\\ y_{i}^{\prime}\sim\mathcal{U}(y_{\min},y_{\max})\\ z_{i}^{\prime}\sim\mathcal{U}(z_{\min},z_{\max})\\ \end{array}\right.$
Upsample (CCS) $P_{new}=P\cup\{p_{i}+\Delta p|\Delta p\sim\mathcal{U}(-0.1,0.1),$
$p_{i}\in RS(P,SEV)\}$
TABLE II: Noise-induced corruptions. SEV (severity) refers to the significance of corruption, whose definition varies based on the type of corruption. C refers to a constant value, whose concrete value varies in different places. RS refers to a random subset function, which samples a subset from the input based on the given $SEV$. $\Delta p$ refers to a point in CCS, so $+\Delta p$ means we change the three coordinates in CCS. $\Delta r$ refers to the range value, $+\Delta r$ means we only change the range value in SCS.
We show the details of noise perturbations in Table II. One frame of point cloud is annotated as $\mathcal{P}=\{p_{i}|i=1,2,...N\}$, where $N$ is the number of points in the point cloud. We apply three distributions of noises (Gaussian, Uniform, and Impulse) to two coordinate systems (Cartesian and Spherical, CCS and SCS for short) to simulate different noise sources. In CCS, one point is represented by $p=\{x,y,z\}$, i.e., Euclidean distance to each planes, while in SCS, one point is represented by $p=\{r,\phi,\theta\}$, where $r$ is the range, i.e., Euclidean distance from the origin to the point, $\phi$ is the azimuthal angle and $\theta$ is the polar angle. Noises in CCS represent inaccuracies caused by the vibration or rotating of the sensor [39] while SCS can better represent inaccurate measurement of time of flight (ToF) [31]. In addition to re-positioning existing LiDAR points, we synthesize new points to simulate environment noises, such as dust in the air, through background noise and upsample applied to CCS. Background noise adds points randomly in the entire space, while upsample adds points close to randomly selected points.
Density Perturbations.
Corruption Formula to Generate Noise
Local Inc $P_{new}=P\cup P_{add},$
$P_{add}=\bigcup_{i=0}^{SEV}IP(NN_{100}(p_{i}))$
Local Dec $P_{new}=P\setminus P_{rm},$
$P_{rm}=\bigcup_{i=0}^{SEV}RS(NN_{100}(p_{i}),C)$
Cutout $P_{new}=P\setminus\bigcup_{i=0}^{SEV}NN_{20}(p_{i})$
Beam Del $P_{new}=RS(P,SEV)$
Layer Del $P_{new}=P\setminus\bigcup_{i\in S}Layer_{i},$
$S=RS(\{1,2,...64\},SEV),P=\bigcup_{i=1}^{64}Layer_{i}$
TABLE III: Density-related corruptions. SEV (severity) refers to the significance of corruption, whose definition varies based on the type of corruption. C refers to a constant value, whose concrete value varies in different places. IP refers to the 3D interpolation function. $NN_{k}$ refers to the k-nearest neighbor function. RS refers to a random subset function, which samples a subset from the input based on the given number.
The density of points can be changed due to internal (e.g., malfunction of the equipment) or external (e.g., occlusion between objects and reflection on object surface) factors [7, 8]. We simulate five density-related corruptions, local density increase, local density decrease, cutout, beam deletion, and layer deletion. The definitions of these corruptions are listed in Table III.
III-B Improving Robustness of LiDAR-Based Pose Estimation
Re-training. Learning-based systems are data-driven and could learn to fit themselves into a data distribution. We expect these systems to learn how to handle the corrupted data by themselves through re-training, which is a common enhancement and defensive strategy for deep learning-based LiDAR-based algorithms [27, 29, 30].
Denoising. Handcraft algorithms are expected to be tuned and adapted by experts once inefficiencies are identified. Therefore such improvements are achieved case by case and lack of generality. Hence in this work, for non-learning-based pose estimation systems, we focus on denoising the data, which is a more general way. Point cloud denoising removes unwanted noises from LiDAR scans to improve their quality. There are three categories of denoising, i.e., filter, optimization, and deep learning-based, each with its advantages and disadvantages [37]. In this work, we adopted a filter-based method, i.e., the bilateral filter by [40], which is reliable, and significantly faster than the alternatives while preserving the intensity with readily available implementation.
IV Experiment Setup
IV-A Subject LiDAR-Based Pose Estimation Systems
TABLE IV: Subject Systems. The errors are reported based on our replication.
Category Dataset Year Method Error
Odometry MULLS [24] 2021 Feature-based 0.56%
KISS-ICP [15] 2023 Direct 0.51%
Delora [19] 2021 Learning-based 7.78%
NeRF-LOAM [22] 2023 NeRF-based 1.70%
Localization LocNDF [21] 2023 NeRF-based 0.059m
We experiment with four SOTA LiDAR odometry and one LiDAR localization systems, as shown in Table IV. We cover one representative work from different types of mainstream point cloud registration methods, i.e., feature-based, direct, and learning-based methods [10]. We also include two neural distance field-based approaches, which have become popular since the success of the NeRF [20].
MULLS [24] is a feature-based LiDAR odometry that uses ground filtering and principal components analysis to extract features from the point cloud. It builds local sub-maps using the previous point cloud and match the current frame with local sub-maps by the proposed multi-metric linear least square iterative closest point algorithm to compute the pose. KISS-ICP [15] is a direct matching LiDAR odometry that combines point-to-point ICP with adaptive thresholding, a robust kernel, universal motion compensation, and point cloud subsampling. Delora [19] is a learning-based, unsupervised LiDAR odometry. It projects the 3D point cloud to 2D image representations and employs ResNet to extract features from the 2D image. Then it predicts the translation and rotation by feeding the features into a neural network. NeRF-LOAM [22] employs a neural network to predict the signed distance field (SDF). It first builds the previous frames into an octree. Then it samples intersected points between laser rays and voxels in the previously built octree. Given sampled points, the neural network predicts the SDF. Since sampled points are transformed to the same coordinate system as the previously built octree by the pose to be estimated, NeRF-LOAM optimizes the pose, neural network, and voxel embedding jointly by minimizing the loss between predicted SDF and real SDF. Different from NeRF-LOAM, LocNDF trains a neural network offline to predict the SDF. Given the online LiDAR scan, it tries to find a pose $P$ so that the neural network can produce minimal SDF when sampled points are transformed by $P$.
IV-B Datasets
We use KITTI Visual Odometry/SLAM Evaluation 2012 (KITTI) [41] to evaluate the four odometry systems, Delora, MULLS, KISS-ICP, and NeRF-LOAM. KITTI 00 to 10 sequences are released with ground truths. We conducted our experiment on sequences 09 and 10 because (1) Delora was trained on sequences 00-08 hence only 09 and 10 remain. (2) sequences 09 and 10 are of moderate sizes which facilitate the large number of perturbations in a reasonable amount of time and computation resources.
Apollo South-Bay [26] contains completed ADS recordings covering various driving scenarios. In this work, we used a subset of ColumbiaPark-3, consisting of 700 testing frames, which is consistent with LocNDF’s pre-trained model [21].
IV-C Evaluation Metrics
For the evaluation of the precision of pose estimation, we adopted the relative pose error (RPE) [41]. RPE is widely used in previous work [10] to measure and compare the performance of pose estimation. RPE is calculated as:
$RPE_{rot}(\mathscr{F})=\frac{1}{|\mathscr{F}|}\sum_{(i,j)\in\mathscr{F}}\angle[(\hat{p}_{j}\ominus\hat{p}_{i})\ominus(p_{j}\ominus p_{i})]$ (2)
$RPE_{trans}(\mathscr{F})=\frac{1}{|\mathscr{F}|}\sum_{(i,j)\in\mathscr{F}}||(\hat{p}_{j}\ominus\hat{p}_{i})\ominus(p_{j}\ominus p_{i})||_{2}$ (3)
where $RPE_{trans}$ measures the relative translation error and $RPE_{rot}$ measures the relative rotation error between the estimated pose and the ground truth, $\mathscr{F}$ is a set of frames $(i,j)$, $\hat{p}$ is the estimated pose and $p$ is the ground truth pose, $\ominus$ is the inverse compositional operator [42] and []delimited-[$\angle$[$\cdot$]∠ [ ⋅ ] is the rotation angle.
V Experiment Results
V-A Evaluating Robustness of LiDAR-Based Pose Estimation
Method. We applied our robustness framework to evaluate the subject odometry systems. We first perturb the KITTI and Apollo-SouthBay datasets with the corruptions in Table I. Note that NeRF-LOAM is computationally expensive (i.e., requiring at least 24GB GPU memory and days to finish one a single sequence), so we took a statistically-significant sample of frames from KITTI for NeRF-LOAM as a proof of concept.
Figure 2: Performance of each system on corrupted KITTI LiDAR Odometry Sequence 09 and 10. Errors with significantly different magnitudes from the majority are aligned at their magnitude and the actual values are labeled in the figure. For example, an error of 12.79% is aligned to “$>$10.0”. If multiple errors overlap at the same magnitude, we label the two highest severity levels.
Results. We show the results of RPEtrans in Figures 2. Due to the space limitation, we do not present the RPErot in the paper as it is correlated with RPEtrans. We share the detailed results of RPErot in our public artifact.
For the non-learning systems, we find that they have degraded performance on almost all kinds of LiDAR perturbed data compared to the baseline (i.e., un-perturbed data). Furthermore, each system is vulnerable to a particular type of perturbation, where their performance drops significantly. The error of MULLS increases from below 0.5% to up to 0.8% on most of the corruptions. However, the error of MULLS with “local density decrease” (severity 5) increased to more than 12% for KITTI 09 and 18% for KITTI 10. Similarly, kiss-icp is especially vulnerable to “background noise”. While the error only increased by less than 0.2% in most cases, the error of kiss-icp caused by background noise increases up to 80% (with severity 5 on KITTI 10) from less than 0.6% RPEtrans. Moreover, even the lightest background noise (severity 1) causes a significant error increase from less than 0.6% to more than 3% RPEtrans on KITTI 10.
We observe different results for the learning-based systems, Delora and NeRF-LOAM. Delora is vulnerable to most perturbations, especially those related to noise. The RPEtrans error increases from less than 10 percent to tens of percent (up to 50%). Lastly, NeRF-LOAM is quite different from others. We find the results differ very much between the two KITTI sequences, indicating that NeRF-LOAM’s robustness results highly depend on data. In KITTI 09, NeRF-LOAM is not sensitive to most perturbations except for one. While in KITTI 10, NeRF-LOAM is shown to be highly sensitive in most cases. Also, interestingly, in many cases, it shows better robustness with a higher severity than a lower severity. For example, for the cutout perturbation on KITTI 10, NeRF-LOAM performs poorly with severity 4, however, the performance is very well with severity 5 which is close to the baseline. NeRF-LOAM optimizes a mini-network online for tens of iterations. We hypothesize that the poor and unstable results are because NeRF-LOAM is far from being optimized in tens of iterations for these cases.
For the LiDAR localization system LocNDF, our evaluation shows that it is very robust against all types of corruption. The robustness is at the cost of significant upfront investments (in terms of multiple scans of the environment to build the LiDAR map) and its inability to work in unexplored environments.
Our results show that the SOTA LiDAR localization system, LocNDF, is robust against all point cloud corruptions. SOTA LiDAR odometry systems are vulnerable to one or multiple types of point cloud corruptions.
V-B Improving Robustness of LiDAR-Based Pose Estimation
Method. We only concern MULLS, kiss-icp, and Delora in this denoising experiment, since LocNDF is robust against all kinds of corruption, while NeRF-LOAM could not produce consistent results between different trials on the same data. Firstly, we apply the bilateral filter to denoise the corrupted data and then re-run all the algorithms on the denoised data. Furthermore, since Delora uses a learning-based approach, we augment Delora’s training set (i.e., KITTI 00 to 08 sequences) by applying the proposed point cloud corruptions and re-train Delora on the augmented data.
We focus on denoising the most representative corruption types that affect the algorithms most, i.e., local density decrease (MULLS and Delora), background noise (kiss-icp), uniform noise (Delora), Gaussian noise (Delora), and rain/snow (moderate effects on all systems). We denoise these types of corruption at severity 1, 3, and 5. We also denoise the clean data to see if the denoising has side effects on clean data.
Similarly, we fine-tune Delora on the corruptions that affect it most significantly, namely, rain, snow, fog, upsample, uniform noise, Gaussian noise, impulse noise, local density decrease, cutout, and layer deletion. We only augmented the training data with these corruptions at severity 5 and fine-tuned Delora for 6 epochs with the default training hyper-parameters.
Results.
Figure 3: (a) Performance on the denoised data. (b) Performance of the retrained model.
We present the performance of the three systems on the denoised data for KITTI 09 in Figure 3 (a). The results on KITTI 10 are similar and available in our replication package. For MULLS, local density decrease (local_dec) at severity 5 is the corruption that affects it most significantly. We find that the errors even increased by 2.43% after denoising. Besides, we find that denoising could reduce the error by 0.26% for Gaussian noise. The changes are insignificant for the remaining cases (less than 0.1%).
For kiss-icp, background noise (bg_noise) is the corruption that affects it most significantly. We find that the denoising can reduce the error by 0.2% (severity 1) to 3.22% (severity 5), that is, the negative effect of the corruption is fully eliminated (we observed the same effects on KITTI 10, where the errors at severity 5 are reduced by 80%). However, the denoising causes the errors on rain and snow to increase by 0.13% to 0.18%, and on local density decrease at severity 5 to increase by 0.28%. The changes are insignificant for the remaining cases.
For Delora, we find that denoising can reduce its errors in most cases including the baseline, by 0.44% to 16.08%. Similar to MULLS and kiss-icp, we find that the errors in local density decrease (local_dec) increased by 2.74%.
Bilateral filter can effectively mitigate noise-induced corruption in general, and fully eliminate the negative effects of background noise. However, it does not help with density-related corruption (local density decrease).
We present the performance of the Delora after fine-turning for KITTI 09 in Figure 3 (b). We find that re-training can bring a much higher improvement than denoising: the errors are reduced by 2.73% to 35.79% for the baseline and all types of corruption. However, for KITTI 10, despite the overall effects being similar, we observed a slight increase in the errors for the baseline and a few corruptions. It is important to note that our fine-tuning was relatively naive: we only augmented KITTI 00 to 08 with the selected corruptions at severity 5 and used the original learning rate. We believe that further improvements could be achieved by a more refined fine-tuning process, such as augmenting the dataset with the corruption at different severities, and adjusting the learning rate and the ratio between clean and corrupted data.
Re-training with the corruption data is a more effective way to improve the robustness of the learning-based LiDAR pose estimation system (Delora).
VI conclusion
In this work, we extensively evaluated the robustness of LiDAR-based pose estimation systems, including odometry and localization, against 18 synthetic real-world point cloud corruptions. Our findings indicate that all SOTA LiDAR-based odometry systems are particularly vulnerable to certain corruption types. We further show that simple bilateral filtering effectively mitigates noise-induced data corruptions, however, density-related corruptions remain challenging. Finally, we show that corruption-based data augmentation not only significantly enhances the robustness of learning-based LiDAR odometry systems but also improves their performance on clean data.