### v3

# Evaluating and Improving the Robustness of LiDAR Odometry and Localization Under Real-World Corruptions

Abstract
LiDAR odometry and localization are two widely used and fundamental applications in robotic and autonomous driving systems. Although state‑of‑the‑art (SOTA) systems achieve high accuracy on clean point clouds, their robustness to corrupted data remains largely unexplored. We present the first comprehensive benchmark to evaluate the robustness of LiDAR pose‑estimation techniques under 18 realistic synthetic corruptions. Our results show that, under these corruptions, odometry position errors escalate from 0.5% to more than 80%, while localization performance stays consistently high. To address this sensitivity, we propose two complementary strategies. First, we design a lightweight detection-and-filter pipeline that classifies the point cloud corruption and applies a corresponding filter (e.g., bilateral filter for noise) to restore the point cloud quality. Our classifier accurately identifies each corruption type, and the filter effectively restores odometry accuracy to near-clean data levels. Second, for learning‑based systems, we show that fine-tuning using the corrupted data substantially improves robustness across all tested corruptions and even boosts performance on clean point clouds on one data sequence.
I Introduction
Reliable pose estimation is critical in robotics and autonomous driving systems (ADS), where even minor errors can lead to catastrophic failures in navigation, planning, and control [1, 2]. LiDAR-based pose estimation is widely adopted for its superior reliability and performance, particularly when Global Navigation Satellite System (GNSS) signals are unreliable (e.g., urban canyons) or unavailable (e.g., indoors).
To this end, extensive research has focused on improving LiDAR pose estimation, i.e., odometry and localization. Particularly, LiDAR odometry estimates an agent’s incremental motion between consecutive LiDAR scans, enabling reliable self-tracking and map building in unknown environments. In contrast, LiDAR localization determines an agent’s absolute pose within a pre-built map, supporting accurate navigation in structured settings such as urban road networks. These methods vary in their registration strategies, including direct scan alignment [3, 4], feature-based matching [5], and, more recently, learning-based approaches [6, 7] to improve generalization across diverse environments.
Despite the success, LiDAR odometry and localization face inherent challenges on robustness caused by non-clean LiDAR data due to real-world corruptions. Prior works reveal that point cloud data (PCD) may be corrupted or noisy, caused by real-world environment challenges, such as adverse weather [8], sensor noises [9] or dusts [10], reflective surfaces [11], and occlusions caused by dynamic environments [12]. This issue is particularly pronounced in learning-based approaches, which often lack transparency and interpretability, making it difficult to understand their decision-making processes and limiting their ability to generalize to scenarios that diverge from the training data [13].
While prior studies have examined specific aspects of robustness, e.g., occlusion in  [12] and security threats via adversarial attacks  [14], these works are limited in scope and do not offer a comprehensive robustness benchmark for LiDAR pose estimation systems. Notably, there is no systematic evaluation of how a wide range of real-world point cloud corruptions affect LiDAR odometry and localization, particularly in learning-based approaches, which are known to be more sensitive to data quality. Furthermore, existing efforts fall short in proposing methods to improve the robustness of LiDAR pose estimation against such corruptions.
In this work, we propose a systematic evaluation framework, RobustLOL, to address the research gap regarding the impact of PCD corruptions on LiDAR odometry and localization, especially learning-based methods which are known to be sensitive to such corruptions. Furthermore, to mitigate these adverse effects, we further introduce and experiment with two complementary strategies to enhance robustness. First, we design a novel and lightweight detection‑and‑filter pipeline and validate its effectiveness on noise‑induced corruptions. Second, for learning‑based methods, we show that fine-tuning on corrupted data significantly enhances robustness. Our key contributions are as follows:
  • We propose RobustLOL, a novel comprehensive robustness benchmark for LiDAR odometry/localization systems.
  • We propose a novel lightweight detection-and-filter method which enables accurate, efficient, generalizable classification and restoration of corrupted PCD.
  • We investigate various mitigation strategies, including detection-and-filter and fine-tuning to defend the LiDAR odometry/localization systems against PCD corruptions.
  • Public release of code and data for reproducibility [^1] https://doi.org/10.5281/zenodo.17128983.
Our framework reveals that SOTA LiDAR odometry is sensitive to at least one type of PCD corruptions. Our detection-and-filter can effectively defend the noise-induced corruptions for non-learning systems and corruption augmented fine-tuning can bring learning-based systems to near-clean level.
II Background and Related Works
II-A LiDAR-Based Odometry and Localization
LiDAR-based odometry and localization allow autonomous agents to track their motions (odometry) and get their absolute position and orientation in a map (localization) in 3D environments. The core of LiDAR-based odometry and localization is point cloud registration, the process of aligning two LiDAR scans to determine their relative poses.
II-A1 Point Cloud Registration
Given a source point cloud $\mathcal{P}\in\mathbb{R}^{3\times N}$ and a target point cloud $\mathcal{T}\in\mathbb{R}^{3\times M}$, point cloud registration solves:
$\mathbf{R}^{*},\mathbf{t}^{*}=\underset{\mathbf{R}\in SO(3),\mathbf{t\in\mathbb{R}^{3}}}{\arg\min}\ \text{dist}\left(\mathbf{R}\mathcal{P}+\mathbf{t},\mathcal{T}\right)$ (1)
where $\mathbf{R}$ is a 3-DoF rotation matrix and $\mathbf{t}$ is a 3-DoF translation vector. The source point cloud is rotated by $\mathbf{R}$ and translated by $\mathbf{t}$. The function $\text{dist}(\cdot,\cdot)$ quantifies the errors between the corresponding points in $\mathbf{R}\mathcal{P}+\mathbf{t}$ and $\mathcal{T}$. In other words, this equation seeks the optimal pose $\mathbf{R}^{*},\mathbf{t}^{*}$ that aligns the source point cloud $\mathcal{P}$ to the target point cloud $\mathcal{T}$.
Various approaches are proposed to solve this problem, evolving from geometry-based matching methods to modern learning-based methods. They can be broadly categorized into direct matching, feature-based, projection-based, and neural distance field-based methods. Direct methods like iterative closest point (ICP) [15, 4] align raw points by iteratively minimizing the distance between nearest neighbor points. Normal distribution transform (NDT) is another classic direct method [3], which partitions the target point cloud into a grid of cells, with each cell modeling the local distribution of points as a Gaussian. Instead of minimizing the distances between points, it optimizes the likelihood measure that evaluates how well the source point cloud fits the normal distribution in the target grid. A recent breakthrough is LiDAR Odometry and Mapping (LOAM) [5] which advances the point cloud registration by extracting robust features (e.g., edges and planar regions) from point clouds and matching the features. With the development of deep learning, researchers further improved feature extraction from hand-crafted features to learned features [6, 7]. The success of Neural Radiance Fields (NeRF) [16] in modeling implicit 3D scenes has inspired the efforts to explore of the application in the LiDAR-based pose estimation [17, 18].
II-A2 LiDAR Odometry
LiDAR odometry tracks real-time motion by aligning consecutive scans. While point cloud registration offers high accuracy, accumulated errors lead to significant long-term drift. To correct this, post-processing techniques like loop closure [19] are often used. SOTA odometry systems cover the various approaches as discussed above [4, 5, 20, 7, 6, 18].
II-A3 LiDAR Localization
LiDAR localization determines the global 6-DoF pose by aligning an online scan with a pre-built map, enabling centimeter-level accuracy critical for fully autonomous driving systems. While some localization systems estimate only 3-DoF position, our work focus on 6-DoF pose estimation. Compared to scan-to-scan registration in odometry, scan-to-map registration in localization is more challenging due to the vast search space. To address this problem, most methods first obtain an initial pose via external signals (e.g., GNSS) [21] or global position recognition [22], then perform the point cloud registration to estimate the fine pose. Although one-stage global pose estimation approaches exist, they generally underperform compared to those with initialization [22]. Similar to odometry, SOTA localization approaches also cover the full spectrum of registration methods [21, 23, 24, 17].
II-A4 Multi-Sensor Fusion Approaches
Multi-sensor fusion (MSF) combines complementary sensors, such as LiDAR, cameras (vision odometry and localization), IMUs (high-frequency motion data), and GNSS (global localization), to enhance pose estimation [21]. In this work, we focus on LiDAR-only pose estimation systems to study LiDAR’s standalone sensitivity to noises in pose estimation.
II-B Robustness of LiDAR-based Systems
Previous research has extensively explored the robustness and security of LiDAR-based 3D obstacle detection. Several studies have also introduced realistic adverse weather simulations to evaluate the impact of environmental conditions on PCD, including rain [25, 26], snow [27], and fog [28]. Researchers have also systematically benchmarked 3D obstacle detection under common point cloud corruptions [29] and investigated its vulnerability to adversarial attacks [30, 31].
For LiDAR pose estimation, recent work has investigated the security of localization and odometry against adversarial attacks [14, 32, 33]. The study most closely related to ours is Laconte et al. [12], which derives a closed-form approximation of the worst-case localization error for the ICP algorithm [15] as a function of the number of corrupted points. Their theoretical bound offers useful insight into how adversarial outliers can degrade a linearized (vanilla) ICP. However, SOTA odometry and localization systems typically employ more advanced ICP variants or alternative registration approaches, and the idealized bounds in [12] do not fully reflect the degradation under realistic corruptions. Our work complements this research gap by performing a systematic empirical evaluation of a broad set of odometry and localization pipelines, including both learning-based and classical registration methods, under a comprehensive suite of realistic corruption scenarios.
III Methodology
In this section, we introduce a unified framework to evaluate LiDAR-based odometry and localization systems under common data corruptions, and we describe two defense strategies: a detection-and-filter pipeline and model fine-tuning on augmented data.
Category Perturbation Type
Weather (5) rain, snow, rain + wet ground (rain_wg),
snow + wet ground (snow_wg), fog
Noise (8) background noise (bg_noise), upsample,
uniform noise in CCS (uni_noise),
gaussian noise in CCS (gau_noise),
impulse noise in CCS (imp_noise),
uniform noise in SCS (uni_noise_rad),
gaussian noise in SCS (gau_noise_rad),
impulse noise in SCS (imp_noise_rad),
Density (5) local density increase (local_inc),
local density decrease (local_dec),
beam deletion (beam_del),
layer deletion (layer_del), cutout
TABLE I: Our robustness framework consists of 18 real-world LiDAR corruptions. CCS refers to Cartesian Coordinate System and SCS refers to Spherical Coordinate System. Their calculation formulas are listed in Table II and III.
III-A RobustLOL: A Robustness Framework for LiDAR Odometry and Localization
Autonomous robots and vehicles are deployed in complex and diverse environments. The performance of LiDAR may degrade under adverse conditions. To rigorously evaluate the robustness of odometry and localization systems, our framework employs 18 different types of PCD corruption (Table I) that simulate adverse weather, internal and external noises, equipment malfunctions, and occlusions.
Weather Perturbations. Adverse weather can negatively affect the quality of LiDAR scans. Randomly distributed airborne droplets from rain, snow, and fog scatter light, corrupting the PCD captured by LiDAR. Furthermore, precipitation results in wet ground that always co-occur with rain or snow, leading to surface reflections that differ significantly from dry ground. Despite the availability of real-world datasets capturing various adverse weather conditions [34, 35], we opt for simulation using clean datasets for three reasons. First, real datasets does not encompass the full spectrum of point-cloud corruptions, whereas simulation enables us to apply every corruptions on the same dataset, ensuring a comprehensive and fair comparison. Second, real recordings are sensor-specific, which can restrict compatibility with certain odometry or localization algorithms [13]; simulation, by contrast, is sensor-agnostic and can be applied to any dataset. Finally, the severity of weather effects in real recordings is fixed and often undocumented, but simulation lets us precisely parameterize corruption intensity and systematically study its impact, leveraging realistic models of fog, rain, and snow [29, 25, 27, 28].
Figure 1: An example of rain and wet ground corruption. The field view and reflection intensities are significantly decayed compared to the clean data. The numbers above the color bar show the intensities of the points with corresponding colors.
Our framework simulates the following weather conditions considering both the direct impact of weather on LiDAR (e.g., droplets causing scattering), as well as impacts of wet grounds on LiDAR (e.g., reflections).
  • Rain or Snow. We use LISA [25, 26], an off-the-shelf rain and snow simulator which use a physics model to simulate the decayed return of laser pulses during rain and snow.
  • Fog. Following previous work [28], we model fog by altering the impulse response, capturing the effect of varying fog densities on the surrounding objects and environment.
  • Rain or Snow + Wet Ground (rain/snow_wg). We combine LISA and Wet Ground Model (WGM) [27] to produce a more realistic simulation than rain/snow alone. WGM builds an optical model to simulate wet ground reflection based on water thickness. We show an example of this corruption in Figure 1.
Noise Perturbations. Noises occur in LiDAR measurement due to internal (e.g., equipment vibration) or external (e.g., dust or strong light) factors [36, 37]. We design perturbations on point cloud to simulate the errors attributed to such factors.
Noise Type Formula to Generate Noise
Gaussian (CCS) $p_{i}+\Delta p*SEV,\;\Delta p\sim\mathcal{N}(0,1)^{3},p_{i}\in P$
Uniform (CCS) $p_{i}+\Delta p*SEV,\;\Delta p\sim\mathcal{U}(-1,1)^{3},p_{i}\in P$
Impulse (CCS) $p_{i}+\Delta p*C,\;\Delta p\sim\mathcal{U}\{-1,1\}^{3},$
$p_{i}\in RS(P,SEV)$
Gaussian (SCS) $p_{i}+\Delta r*SEV,\;\Delta r\sim\mathcal{N}(0,1),p_{i}\in P$
Uniform (SCS) $p_{i}+\Delta r*SEV,\;\Delta r\sim\mathcal{U}(-1,1),p_{i}\in P$
Impulse (SCS) $p_{i}+\Delta r*C,\;\Delta r\sim\mathcal{U}\{-1,1\},$
$p_{i}\in RS(P,SEV)$
Background $P_{new}=P\cup\{p_{i}^{\prime}|j=1,2,...,SEV\}$
[…]
Upsample $P_{new}=P\cup\{p_{i}+\Delta p|\Delta p\sim\mathcal{U}(-0.1,0.1),$
(CCS) $p_{i}\in RS(P,SEV)\}$
TABLE II: Noise-induced corruptions. SEV (severity) is the significance of corruption, whose definition varies based on the type of corruption. C is a constant value, whose concrete value varies in different scenarios. RS is a random subset function, which samples a subset from the input based on the given SEVSEV. $\Delta p$ is a point in CCS, so $\Delta p$ means we change the three coordinates in CCS. $\Delta r$ refers to the range value, $\Delta r$ means we only change the range value in SCS.
We show the details of noise perturbations in Table II. One frame of point cloud is annotated as $\mathcal{P}=\{p_{i}|i=1,2,...N\}$, where $N$ is the number of points in the point cloud. We apply three distributions of noises (Gaussian, Uniform, and Impulse) to two coordinate systems (Cartesian and Spherical, CCS and SCS for short) to simulate different noise sources. In CCS, one point is represented by $p=\{x,y,z\}$, i.e., Euclidean distance to each planes, while in SCS, one point is represented by $p=\{r,\phi,\theta\}$, where $r$ is the range, i.e., Euclidean distance from the origin to the point, $\phi$ is the azimuthal angle and $\theta$ is the polar angle. Noises in CCS represent inaccuracies caused by the vibration or rotating of the sensor [38] while SCS can better represent inaccurate measurement of time of flight (ToF) [29]. In addition to re-positioning existing LiDAR points, we synthesize new points to simulate environment noises, such as dust in the air or strong external light, through background noise and upsample applied to CCS. Background noise adds points randomly in the entire space, while upsample adds points close to randomly selected points.
Density Perturbations.
Corruption Formula to Generate Noise
Local Inc $P_{new}=P\cup P_{add},$
$P_{add}=\bigcup_{i=0}^{SEV}IP(NN_{100}(p_{i}))$
Local Dec $P_{new}=P\setminus P_{rm},$
$P_{rm}=\bigcup_{i=0}^{SEV}RS(NN_{100}(p_{i}),C)$
Cutout $P_{new}=P\setminus\bigcup_{i=0}^{SEV}NN_{20}(p_{i})$
Beam Del $P_{new}=RS(P,SEV)$
Layer Del $P_{new}=P\setminus\bigcup_{i\in S}Layer_{i},$
$S=RS(\{1,2,...64\},SEV),P=\bigcup_{i=1}^{64}Layer_{i}$
TABLE III: Density-related corruptions. SEV (severity) refers to the significance of corruption, whose definition varies based on the type of corruption. C refers to a constant value, whose concrete value varies in different places. IP refers to the 3D interpolation function. $NN_{k}$ refers to the k-nearest neighbor function. RS refers to a random subset function, which samples a subset from the input based on the given number.
The density of points can be changed due to internal (e.g., malfunction of the equipment) or external (e.g., occlusion between objects and reflection on object surface) factors [11, 39]. We simulate five density-related corruptions, local density increase, local density decrease, cutout, beam deletion, and layer deletion. The definitions of these corruptions are listed in Table III.
III-B Improving the Robustness of LiDAR Odometry and Localization
We propose and experiment with two types of approaches to improve the robustness of such LiDAR systems.
III-B1 For All LiDAR Odometry and Localization.
We propose a novel detection-and-filter pipeline to improve the robustness of LiDAR odometry and localization systems. This pipeline is algorithm-agnostic and can be widely applied to algorithms of LiDAR pose estimations. Our novel pipeline consists of two stages: (1) corruption detection via a DNN-based classifier, and (2) targeted corruption filtering based on the corruption type.
Stage 1 – Corruption Detection. As odometry and localization are real-time applications, we focus on techniques that are inexpensive in terms of time and memory cost. Hence, we first convert each 3D LiDAR scan into a 2D range image via spherical projection similar to [7], creating a $64\times 1024$ grid, where each pixel stores the depth and intensity of the point in that cell closest to the agent. This representation preserves the spatial structure of the scene while enabling efficient convolutional processing.
Using these range images, we train a lightweight CNN (Figure 2) to classify data corruptions (e.g., rain, noises) in LiDAR data on our augmented dataset by minimizing categorical cross-entropy loss. This loss guides the model to improve classification accuracy across different corruption types through successive training epochs.
The CNN consists of three convolutional layers with increasing feature sizes (32, 64, and 128), each followed by batch normalization and ReLU activation. A final adaptive average pooling layer reduces the output to a 128-dimensional vector to produce class scores, where the class with highest probability is selected as the predicted class for the filtering stage. This lightweight architecture is designed for real-time or embedded deployment, balancing classification accuracy with minimal computational overhead.
Figure 2: Architecture of CNN for corruption classification.
Stage 2 – Corruption Removal. After identifying the corruption type, we apply the corresponding filter to mitigate its effects and restore point-cloud fidelity. While our framework supports any paired detector–filter module (e.g., precipitation or density filters), in this work we concentrate on noise-induced corruptions because other point-cloud-based systems are sensitive to noises [29] and denoising methods are mature and effective in removing noises. Thus, once the classifier flags a scan as noise-corrupted, we pass it to a bilateral filter [40] for denoising. This modular design allows additional filters to be plugged in for other corruption categories.
III-B2 For Learning-Based LiDAR Odometry and Localization.
Learning-based approaches inherently adapt to the characteristics of their training data. Hence, to defend these systems against the data corruptions, we augment the training set with perturbations shown to degrade performance by our evaluation framework and then fine-tune the pretrained models on this augmented dataset. We fine-tune the model using the original training recipe from the released checkpoint until it converges on the augmented dataset. This strategy encourages the neural networks to learn invariant features directly from the data distribution, and has been shown to enhance resilience in LiDAR-based systems [25, 27, 28].
IV Experiment Setup
IV-A Subject LiDAR Pose Estimation Systems
Category Name Year Method
Odometry MULLS 2021 Feature-based
KISS-ICP 2023 Direct
DeLORA 2021 Learning-based
NeRF-LOAM 2023 Learning-based
Localization LocNDF 2023 Learning-based
TABLE IV: Subject LiDAR odometry and localization techniques.
We evaluate four odometry and one localization systems, as shown in Table IV. The subject systems comprise two representative classical methods, MULLS and kiss-icp, and three learning-based approaches, DeLORA, NeRF-LOAM, and LocNDF. MULLS [20] is a feature-based LiDAR odometry that uses ground filtering and principal components analysis to extract features from the point cloud. It builds local sub-maps using the previous point cloud and match the current frame with local sub-maps by the proposed multi-metric linear least square iterative closest point algorithm to compute the pose. KISS-ICP [4] is a direct matching LiDAR odometry that combines point-to-point ICP with adaptive thresholding, a robust kernel, universal motion compensation, and point cloud subsampling. DeLORA [7] is a learning-based, unsupervised LiDAR odometry. It projects the 3D point cloud to 2D image representations and employs ResNet to extract features from the 2D image. Then it predicts the translation and rotation by feeding the features into a neural network. NeRF-LOAM [18] employs a neural network to predict the signed distance field (SDF). It first builds the previous frames into an octree. Then it samples intersected points between laser rays and voxels in the previously built octree. Given sampled points, the neural network predicts the SDF. Since sampled points are transformed to the same coordinate system as the previously built octree by the pose to be estimated, NeRF-LOAM optimizes the pose, neural network, and voxel embedding jointly by minimizing the loss between predicted SDF and real SDF. Different from NeRF-LOAM, LocNDF trains a neural network offline to predict the SDF. Given the online LiDAR scan, it tries to find a pose $P$ so that the neural network can produce minimal SDF when sampled points are transformed by $P$.
IV-B Datasets
We use KITTI Visual Odometry/SLAM Evaluation 2012 (KITTI) [41] to evaluate the four odometry systems, DeLORA, MULLS, KISS-ICP, and NeRF-LOAM. The reason is that KITTI is one of the most popular dataset for odometry and all of the selected odometry systems have been evaluated on KITTI. KITTI 00 to 10 sequences are released with ground truths. We conducted our experiment on sequences 09 and 10 because (1) DeLORA was trained on sequences 00-08 hence only 09 and 10 remain. (2) sequences 09 and 10 are of moderate sizes which facilitate the large number of perturbations in a reasonable amount of time and computation resources.
We use Apollo South-Bay [24] to evaluate LocNDF. Apollo South-Bay contains completed ADS recordings covering various driving scenarios. LocNDF released the pretrained model and pre-built map on this dataset. In this work, we used a subset of ColumbiaPark-3, consisting of 700 testing frames, which is consistent with LocNDF’s pre-trained model [17].
IV-C Evaluation Metrics
For the evaluation of the precision of pose estimation, we adopted the translational relative pose error (RPEtrans[41]. RPEtrans is widely used in previous work [13] to measure and compare the performance of pose estimation. RPEtrans is calculated as:
[…]
where $\mathscr{F}$ is a set of frames $(i,j)$, $\hat{p}$ is the estimated pose and $p$ is the ground truth pose, $\ominus$ is the inverse compositional operator [42]. A pose $p$ consists of a rotation matrix $R$ and a translation vector $t$.
V Experiment Results
V-A RobustLOL: Robustness Results of LiDAR Odometry and Localization
Method. We applied our robustness framework to evaluate the subject odometry and localization systems. We first perturb the KITTI and Apollo-SouthBay datasets with the corruptions in Table I. Note that NeRF-LOAM is computationally expensive (i.e., requiring at least 24 GB GPU memory and several days to finish a single sequence), so we took a statistically-significant sample of frames from KITTI for NeRF-LOAM as a proof of concept.
Figure 3: Performance of each system on KITTI 09.
Figure 4: Performance of each system on KITTI 10.
Results. Figure 3 and 4 present RPEtrans on KITTI 09 and 10 for all corruptions of severities 1, 3, and 5. We omit the results of LocNDF as we observe no performance degradation under any corruption in our experiment. We share the complete results in the released artifacts.
Non-learning systems show overall robustness to most perturbations with each system particularly sensitive to certain corruption types, leading to significant performance drops. MULLS is sensitive to local density decrease (local dec) at severity 5 where RPEtrans increases from 0.3% to over 18% and kiss-icp is sensitive to background noise where RPEtrans rises sharply from less than 0.6% on the clean data to over 80%. Our investigation shows that local density reduction disrupts key point clusters which MULLS relies on, making its feature extraction unreliable. On the other hand, background noise introduces uniformly distributed noisy points which disrupt kiss-icp’s direct point-on-point registration.
In contrast, learning-based odometry systems, i.e., DeLORA and NeRF-LOAM are more sensitive to most corruptions. Notable exceptions include background noise, spherical-coordinate noises (uniform, Gaussian and impulse noises (SCS)), local density increase, and beam deletion, where performance remains relatively stable due to DeLORA’s 2D range image representation. During projection, multiple 3D points may map to the same pixel, with only the closest point retained, which significantly limit the effect of these perturbations. Thus, background noise appears as mild pixel-level noise, spherical-coordinate noise alters depth without affecting pixel locations, beam deletion has limited impact due to point cloud redundancy, and local density increase adds points that often overlap with existing pixels, preserving image structure. In contrast, other corruptions introduce geometric distortions or remove key features, leading to more severe performance drops. Finally, we find that NeRF-LOAM shows substantial instability, particularly on KITTI 10. In some cases, increasing corruption severity produces lower errors which is counterintuitive. To investigate this, we repeated the experiments under each corruption two additional times (three runs in total) with identical inputs. The resulting trajectories and final errors vary substantially across runs. We attribute this variability to random network initialization and the limited number of online optimization iterations.
Our systematic evaluation reveals key vulnerabilities in various LiDAR odometry systems.
V-B Detection-and-Filter Pipeline
Figure 5: (a) $RPE_{trans}$ with our detection-and-filter pipeline. (b) $RPE_{trans}$ of the fine-tuned DeLORA on the augmented dataset. Colored bars (light gray, medium gray, dark gray) report the RPE at severity levels 1, 3, and 5 after applying our pipeline or fine-tuning; the hatched white bars with “/” show the corresponding errors without our pipeline or fine-tuning.
Method. We apply the detection-and-filter pipeline on all systems except NeRF-LOAM because it could not produce consistent results between different trials on the exact same data and LocNDF because it is robust against all proposed corruptions.
We implement the classifier proposed in Section III-B1 using PyTorch 1.13.1+cu117 and train it to recognize eight representative corruptions: rain, snow, fog, uniform noise (CCS), Gaussian noise (CCS), background noise, local density increase, and local density decrease. Due to their similar scattering effects, we merge rain and snow into one class; likewise, uniform and Gaussian noise (CCS) are grouped together for their comparable distortion patterns. We train the model on KITTI 09 (80% train, 20% validation) for 10 epochs using a batch size of 8, a learning rate of 0.001 with the Adam optimizer. We apply it on KITTI 10 and we use the bilateral filter to restore the noise-corrupted PCD. All experiments are run on Ubuntu 22.04 LTS with an AMD Ryzen 32-core CPU, 256 GB RAM, and an NVIDIA GeForce RTX 3090 GPU.
Results. Our classification model achieves an average accuracy of 95.54%, as shown in Table V. We notice that clean and fog have a comparably lower accuracy because they are often misclassified as local density increase. The reason is that (i) fog has sparse reflections which increase the points density, and (ii) occasional high-reflecting surfaces in clean data also increases points density.
Weather Noise Density
Clean Fog R+S U+G Back Inc Dec
Acc 90.02% 80.83% 98.32% 100% 100% 100% 97.84%
TABLE V: Classification results for seven class of noises, i.e., Clean, Fog, Rain+Snow (R+S), Uniform+Gaussian Noise (U+G), Background Noise (Back), Local Density Increase (Inc), and Local Density Decrease (Dec)
We present the RPEtrans of the three systems incorporated with our pipeline on KITTI 10 in Figure 5 (a). MULLS is shown to be robust against most noise in the previous section. Applying our detection-and-filter pipeline to the noisy data causes negligible performance degradation (less than 0.1%). kiss-icp significantly benefits from the detection-and-filter pipeline, with post-denoising errors returning to the clean-data baseline. DeLORA also has a significant error reduction after filtering noises in data, however, the errors are still much higher than the baseline.
Our classifier could classify noise types accurately and our detection-and-filter pipeline could significantly reduce the error on noise-corrupted point cloud data.
V-C Fine-Tuning
Method. We fine-tune the learning-based system, i.e., DeLORA, on the dataset augmented with the corruptions that affect it most significantly, namely, rain, snow, fog, upsample, uniform noise (CCS), Gaussian noise (CCS), impulse noise (CCS), local density decrease, cutout, and layer deletion. We exclude NeRF-LOAM and LocNDF for the same reasons mentioned in Section V-B.
Results. We present the performance of the fine-tuned DeLORA on corrupted KITTI 10 scans (KITTI 09 results are available in the replication package) in Figure 5 (b). We find that the errors are reduced by up to 35.79% for all types of corruption after fine-tuning which brings the performance close to the clean data. On the baseline, the fine-tuned DeLORA achieves a 2.7% decrease in RPEtrans on KITTI 09 but a 3.3% increase on KITTI 10 compared to the original model. It is important to note that our fine-tuning was comparably simple: we only augmented KITTI 00 to 08 with the selected corruptions at severity 5 and used the original learning rate. We believe that further improvements could be achieved by a more refined fine-tuning process, such as augmenting the dataset with the corruption at different severities, and adjusting the learning rate and the ratio between clean and corrupted data.
Fine-tuning with the corruption data is an effective way to improve the robustness of the learning-based LiDAR odometry system (DeLORA).
VI Conclusion
In this work, we propose RobustLOL, a systematic framework for evaluating the robustness of LiDAR-based odometry and localization systems against 18 types of synthetic real-world point cloud corruptions. Our evaluation reveals critical vulnerabilities in different types of odometry systems, with performance degradation of up to 80%. We further demonstrate that our simple detection-and-filter pipeline can effectively mitigate noise-induced corruptions, restoring performance to near-clean levels. Finally, we show that corruption-based data augmentation substantially enhances the robustness of learning-based systems against point cloud corruptions. Notably, it also improved clean-data performance on KITTI 09, though a slight decrease was observed on KITTI 10.